# Occupancy-aware Trajectory Planning for Autonomous Valet Parking in Uncertain Dynamic Environments

### v1

@@ Abstract
Accurately reasoning about future parking spot availability and integrated planning is critical for enabling safe and efficient autonomous valet parking in dynamic, uncertain environments. Unlike existing methods that rely solely on instantaneous observations or static assumptions, we present an approach that predicts future parking spot occupancy by explicitly distinguishing between initially vacant and occupied spots, and by leveraging the predicted motion of dynamic agents. We introduce a probabilistic spot occupancy estimator that incorporates partial and noisy observations within a limited Field-of-View (FoV) model and accounts for the evolving uncertainty of unobserved regions. Coupled with this, we design a strategy planner that adaptively balances goal-directed parking maneuvers with exploratory navigation based on information gain, and intelligently incorporates wait-and-go behaviors at promising spots. Through randomized simulations emulating large parking lots, we demonstrate that our framework significantly improves parking efficiency, safety margins, and trajectory smoothness compared to existing approaches.
@@ I INTRODUCTION
In busy public spaces, parking consumes a lot of time, fuel and curb space, as drivers loop through parking lots searching for parking, interact with other drivers and pedestrians and perform tight maneuvers into a parking spot. Parking assistant systems that steer a car into a user-selected parking spot are now commercially available [1, 2], however a Type-1 autonomous valet parking system [3] that does not rely on infrastructure support is not yet available. We address the challenge of jointly selecting a parking spot and planning collision-free paths using only onboard sensor information.
Prior research in autonomous parking often has simplified assumptions that undermine realistic applicability. For instance, existing trajectory planners assume a predefined reference trajectory and focus only on static obstacle avoidance [4] or dynamic obstacle avoidance only for two-vehicle scenarios [5], neglecting the problem of identifying an optimal spot dynamically. Moreover, occupancy prediction methods traditionally rely on basic Bayesian filtering techniques without considering the influence of dynamic agents [6]. Their observation models assume constant sensor confidence irrespective of distance from the ego vehicle, disregarding the practical degradation of observation reliability with increased range. Additionally, exploration strategies in previous works [7, 6] overlook the downstream navigation task of executing the actual parking maneuver. In particular, [7] employs a simplistic entropy-based model to represent spot occupancy distributions, limiting its ability for real-time decision-making under uncertainty. Furthermore, the frameworks in [8] assume complete observability of the parking environment, with full Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I) communication capabilities, and restrict themselves to offline spot assignment tasks without real-time strategic adjustments.
@@
To address these gaps, we propose an occupancy-aware trajectory planning framework tailored for uncertain, dynamic parking scenarios with the following contributions:
  • Partial Observation model: To consider partial observability, we introduce a shifted‑rectangle Field-of-View model whose observation confidence decays with distance, thereby modeling the decreasing confidence of spots being available further away.
@@
  • Spot Occupancy estimator: We propose a Bayes filter based approach that predicts the future occupancy of parking spots using the behavior of dynamic agents in the scene. We also model distinct arrival and departure processes for initially vacant and initially occupied spots respectively, to reason about the future occupancy.
  • Strategy Planner: We couple the occupancy beliefs with a cost‑based policy that plans a trajectory for the ego to explore the lot or commit to a spot, balancing efficiency and safety.
  • Waiting for a spot: We enable the ego vehicle to pause briefly near promising spots, capturing imminent vacancies that would otherwise be missed and reducing unnecessary exploration.
We evaluate our system in a simulated parking lot and show that our system reduces time and distance traveled to park and increases safety margins relative to existing methods, confirming that foresighted occupancy reasoning improves the overall valet parking performance.
@@ II Problem formulation and Approach
We consider a parking lot environment, where the autonomous vehicle has prior access to the static map. The vehicle is equipped with onboard sensors that provide partial observations of parking spot occupancy within a limited FoV and also provides motion predictions of surrounding dynamic agents, such as vehicles and pedestrians. The task is to plan trajectories that guide the ego-vehicle to explore the parking lot, find an available parking spot, and park in a safe and efficient manner.
@@ II-A Parking Lot Environment
Consider the parking lot in Fig. 1 where there is a grid of $N_{p}$ parking spots. The centers of all the spots in the lot are given by the set $\mathcal{P}$, where each $s_{i}\in\mathcal{P}$ is the two-dimensional Cartesian coordinates of the center of spot $i$, where $i\in\{1,2,\ldots,N_{p}\}$. We assume that the ego has prior access to the static map, which consists of the parking lot bounds, the centers of the spots $\mathcal{P}$ and information about road lanes (center of road and lane width). The parking lot bounds are modeled as static obstacles as described in [9].
Figure 1: Autonomous ego vehicle (green) with limited FoV (blue/orange) navigating a parking lot with dynamic vehicles (red rectangles) and a pedestrian (red circle). Spots are fully observed inside blue rectangle ($\epsilon=1$), unobserved outside orange ($\gamma=1.5$), and partially observed in between. Static vehicles are in black, and parking spot centers are grey dots.
@@ II-B Vehicle Model
Let $x=(X,Y,\theta)$ be the state of the vehicle, where $(X,Y)$ is the center of the rear axis and $\theta$ is the heading angle. We employ the kinematic bicycle model for vehicle dynamics which is well suited for vehicle at low speeds [10], given as
@@ II-C Observation Model
The limited on-board sensing capabilities is modeled using a FoV observation model based on a “scaled and shifted” infinity norm, with the sensing region biased towards the front of the vehicle.
@@
where $r_{x}$ and $r_{y}$ define the half-lengths of the FoV rectangle along the longitudinal and lateral directions, respectively, while $\zeta$ is the forward shift in ego frame to bias the sensing region ahead of the vehicle. The rotation matrix $R(\theta)$ transforms vectors from the ego frame to the world frame.
@@
Let $O_{t}^{s}\in\{0,1\}$ be the ground-truth occupancy and $Z_{t}^{s}\in\{0,1\}$ be the observed occupancy of spot $s$ at time $t$, where $1:=$ occupied and $0:=$ vacant. The observation accuracy is modeled as the likelihood of correctly observing spot $s$ using the below distance-dependent probability
$p_{c}(d_{(\zeta,r)}(x,s))=P(Z_{t}^{s}=1|O_{t}^{s}=1)=P(Z_{t}^{s}=0|O_{t}^{s}=0),$
$p_{c}(d_{(\zeta,r)}(x,s))=e^{-\alpha\cdot\sigma(d_{(\zeta,r)}(x,s))}\ \textrm{s.t.}\ \alpha,$
$\sigma(d_{(\zeta,r)}(x,s))=\frac{\ln(2)}{\alpha}\left(1+e^{-\alpha_{c}\left(d_{(\zeta,r)}(x,s)-\left(\frac{\gamma-\epsilon}{2}+\epsilon\right)\right)}\right)^{-1}.$ (4)
The function $\sigma(d)$ is a scaled sigmoid centered at $\tfrac{\gamma-\epsilon}{2}+\epsilon$ where $\alpha_{c}\gg 0$, with $\sigma(d)\approx 0$ when $d\approx\epsilon$ and $\sigma(d)\approx\frac{\ln(2)}{\alpha}$ when $d\approx\gamma$. Hence, $p_{c}(d)\to 1$ as $d\to\epsilon$ and $p_{c}(d)\to 0.5$ as $d\to\gamma$. Thus, our model captures how observation confidence decays with distance and varies by direction, unlike prior work [6] that assumes fixed uncertainty.
@@
Given the state $x_{t}\in\mathbb{R}^{3}$ of the ego vehicle at time $t$, parking spot observations $Z_{t}^{s}$ as defined in Section II-C, predictions of $D$ dynamic agents $\{\{\hat{y}^{j}(k)\}_{k=t}^{t+T}\}_{j=1}^{D}$ over a horizon $T$, where $\hat{y}^{j}(k)\in\mathbb{R}^{3}$ is the position and heading of agent $j$ if it is a vehicle, and $\hat{y}^{j}(k)\in\mathbb{R}^{2}$ is the position if agent $j$ is a pedestrian; compute a sequence of control inputs $\{u_{t},u_{t+1},\ldots,u_{t+T}\}$ for the kinematic model (1) that steers the vehicle to a feasible parking spot or explores the lot while avoiding both static obstacles and dynamic agents.
@@ II-D Proposed Approach
Fig. 2 illustrates our approach. At each planning step, we predict future parking spot occupancy over a horizon $T$ using a recursive belief estimator that fuses our FoV observation model (Section II-C) with dynamic agent predictions. Our strategy planner then selects one of three actions—park immediately, wait near a potentially vacant spot, or explore the lot—by minimizing a cost balancing efficiency and safety. Hybrid A* is used in our path planner that generates a smooth, collision-free trajectory for execution, after which the loop repeats with updated observations.
@@
III Spot Occupancy Estimator
At each planning step $t$, the spot occupancy estimator predicts the probability of each parking spot being occupied over a horizon $T$, combining current sensor observations with the predicted trajectories of dynamic agents. This allows the ego vehicle to reason about which spots are vacant in the immediate future, guiding navigation based on anticipated occupancy rather than instantaneous observations.
We estimate parking spot occupancy probabilities using the Bayes Filter predict and update framework [11]. The prediction step models dynamic agents and their interactions with the environment, while the update step incorporates FoV observations as described in Section II-C.
@@ III-A Effect of Dynamic Agents
Each agent $j$’s predicted motion is modeled as a time-indexed Gaussian
@@
capturing both the nominal trajectory $\hat{y}^{j}(k)$ and increasing uncertainty over time. These uncertain trajectories are used to compute the probability that a parking spot $s\in\mathcal{P}$ is occupied by an agent $j$ at time $k$:
@@
where the two terms in (6) are
$E_{d}=\big{(}(\alpha_{d}+1)(\alpha_{d}+e^{\alpha_{1}m_{d}(y^{j}(k),s)})\big{)}^{-1},$ (7)
which is an exponential decay over the Mahalanobis distance [12] between the agent’s position and the spot:
$m_{d}(y^{j}(k),s)=(\hat{y}^{j}(k)-s)^{T}\Sigma(k)^{-1}(\hat{y}^{j}(k)-s),$
and an alignment factor
@@
capturing whether the agent is moving towards the spot, using the predicted velocity $\hat{v}^{j}(k)$ and velocity uncertainty $\sigma_{v}(k)$.
Uncertainty Propagation: We propagate the positional covariance at each timestep using the discrete-time model
@@
where $R(\theta(k))$ is the rotation matrix parameterized by the vehicle’s heading $\theta(k)$ and $Q$ is a diagonal matrix representing the process noise in ego frame. To estimate the uncertainty in velocity, we use the finite difference $v^{j}(k)=\frac{y^{j}(k+1)-y^{j}(k)}{\Delta t}$} and (9).
From (6), the probability that at least one of the $D$ dynamic agents will occupy spot $s$ at time $k$ is
@@
At each future step $k\in\{t,\dots,t+T-1\}$, we predict the prior belief $\bar{b}_{k}^{s}=P(O_{k}^{s}=1)$ based on the behavior of other agents, and update it using noisy, partial observations $Z_{k}^{s}$. The resulting posterior belief is $b_{k}^{s}=P(O_{k}^{s}=1\mid Z_{k}^{s})$.
@@ III-B Prediction
In prior work [6], belief is propagated using a transition model that ignores interactions between dynamic agents and parking spots. In contrast, we explicitly incorporate dynamic agent predictions and model the asymmetry between spots that are initially vacant versus occupied. The initial state of a spot is determined from its observation $Z_{t}^{s}$ under the FoV model in Section II-C. For spots outside the FoV, neither occupancy observations nor agent interactions are available.
@@
    where $q(s,k+1)=P(O_{k+1}^{s}=1|O_{k}^{s}=0)$ is computed from (10) and $\mu_{1}=P(O_{k+1}^{s}=1|O_{k}^{s}=1)$. The term $\mu_{1}$ models the departure probability of a parked vehicle using an exponential distribution, reflecting the memory-less nature of departures. If $\lambda_{d}$ denotes the average departure rate of a parked vehicle, the probability of departure within a time interval $\Delta t$ is
@@
    where $1-q(s,k+1)=P(O_{k+1}^{s}=0|O_{k}^{s}=1)$ is computed from (10) and $\mu_{2}=P(O_{k+1}^{s}=1|O_{k}^{s}=0)$. We model the arrival probability of a vacant spot using a Poisson distribution. If $\lambda_{a}$ denotes the average number of vehicles that arrive with the time interval $\Delta t$, the probability that no vehicle parks in a vacant spot is
@@
    For unobservable spots, the prediction equation relies only on the average departure and arrival rates:
@@
The predicted prior $\bar{b}_{k+1}^{s}$ that is generated from the spot’s occupancy history and dynamic agent predictions serves as the input for the update step to fuse the observations.

### v2

@@ Abstract
Autonomous Valet Parking (AVP) requires planning under partial observability, where parking spot availability evolves as dynamic agents enter and exit spots. Existing approaches either rely only on instantaneous spot availability or make static assumptions, thereby limiting foresight and adaptability. We propose an approach that estimates probability of future spot occupancy by distinguishing initially vacant and occupied spots while leveraging nearby dynamic agent motion. We propose a probabilistic estimator that integrates partial, noisy observations from a limited Field-of-View, with the evolving uncertainty of unobserved spots. Coupled with the estimator, we design a strategy planner that balances goal-directed parking maneuvers with exploratory navigation based on information gain, and incorporates wait-and-go behaviors at promising spots. Through randomized simulations emulating large parking lots, we demonstrate that our framework significantly improves parking efficiency and trajectory smoothness over existing approaches, while maintaining safety margins.
@@ I INTRODUCTION
In busy public spaces, parking consumes significant time, space, and fuel [1, 2]. Drivers often loop through lots such as shown in Fig. 1, searching for spots, interacting with other drivers and pedestrians, and performing tight maneuvers. While existing park assist systems can steer a car into a user-selected spot [3, 4], a Type-1 autonomous valet parking system [5] that operates without infrastructure support remains unavailable. To this end, we propose a framework for Type-1 AVP that integrates an observation model tailored for onboard sensing with unified spot selection and trajectory planning.
Figure 1: Example of a crowded shopping mall parking lot [6].
Several trajectory planners for autonomous parking have been proposed using optimization and learning-based approaches. Some optimization-based methods often focus on static obstacle avoidance [7, 8], while dynamic obstacle avoidance approaches [9, 10] typically rely on predefined reference paths. Time-optimal formulations use dynamic optimization [11, 12] to reduce parking time but require high computational effort, whereas heuristic planners that use Reeds–Shepp curves [13] trade optimality for faster real-world execution. Learning-based approaches leverage deep neural networks [14, 15, 16] and reinforcement learning [17, 18] to generate diverse trajectories in highly constrained and partially unknown environments, but fail to guarantee safety. While these approaches demonstrate effective maneuver execution, they generally assume predefined spot availability and decouple spot selection from trajectory planning.
Methods that focus on parking spot selection often make simplifying assumptions that limit practical applicability. Some approaches ignore spot occupancy prediction [19], while others assume full connectivity between vehicles [20]. Reinforcement learning strategies [21, 22] optimize parking space allocation and reduce search time, but struggle to generalize and guarantee real-time safety. Vanilla Bayesian filtering methods [23] typically neglect the influence of dynamic agents on spot availability, and their observation models [23, 19] assume fixed sensor confidence within the Field-of-View (FoV), overlooking the practical degradation of perception reliability with distance. Exploration-driven strategies [24, 23] focus on information gain but often disregard the downstream task of actual parking. For instance, [24] uses a simplified entropy-based occupancy model that ignores vehicle dynamics, limiting its suitability in dynamic environments. Connected-vehicle frameworks [20, 25] assume Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I) communication, which are effective for offline spot allocation, but do not address the real-time strategic adjustments required in infrastructure-free scenarios.
@@
To address these gaps, we propose an occupancy-aware trajectory planning framework tailored for Type-1 AVP:
  • Distance-aware partial FoV model: We model partial observability by introducing a distance-dependent FoV model that maps range to observation confidence.
@@
  • Spot occupancy estimator: We propose a Bayes filter based approach that predicts the future occupancy of parking spots using the behavior of dynamic agents. We also model distinct arrival and departure equations for initially vacant and occupied spots, respectively, to reason about the future occupancy.
  • Strategy planner: We integrate future spot occupancy probabilities into a cost-based policy that plans the ego vehicle’s trajectory to either explore the parking lot or commit to a spot. The policy allows the ego to pause near promising spots by incorporating a waiting-time cost, enabling it to capture imminent vacancies while maintaining a balance between efficiency and safety.
We evaluate our approach in a simulated parking lot, combining foresighted occupancy reasoning and strategy planning. Our method reduces time and distance traveled while generating smoother and safer paths compared to existing spot selection methods and trajectory planners [23, 20].
@@ II Problem formulation and Approach
We consider an ego vehicle navigating a parking lot, equipped with onboard sensors that provide observations of spot occupancy within a limited FoV, along with motion predictions of dynamic agents such as vehicles and pedestrians. The task for the ego is to plan trajectories that efficiently explore the lot, identify an available spot, and park safely.
@@ II-A Parking Lot Environment
Consider the parking lot in Fig. 2 with $N_{p}$ spots, whose centers are given by $\mathcal{P}=\{s_{1},\dots,s_{N_{p}}\}$, where $s_{i}$ denotes the 2D coordinates of spot $i$. The ego vehicle has prior access to a static map of the lot (boundaries and lane centerlines/widths), but spot occupancy is unknown and must be observed in real time. Parking lot boundaries are treated as static obstacles as given in [26].
Figure 2: Autonomous ego vehicle (green) with limited field of view (blue/orange) navigating an environment simulating a mall parking lot, with the entrance at the top. Predicted motions of dynamic vehicles are shown as red rectangles, and a pedestrian is indicated by a red circle. Spots are fully observed inside blue rectangle ($\epsilon=1$), unobserved outside orange ($\gamma=1.5$), and partially observed in between. Static vehicles are black, and parking spot centers are grey dots.
@@ II-B Vehicle Model
Let $x=[X,Y,\theta]^{\top}$ be the state of the vehicle, where $(X,Y)$ is the center of the rear axis and $\theta$ is the heading angle. We model vehicle dynamics using the kinematic bicycle model [27], which is well suited for low-speed maneuvers:
@@ II-C Observation Model
The vehicle’s limited onboard sensing is modeled via a “scaled and shifted” infinity norm defining the FoV, though other shapes (e.g., ellipses or circles) can be modeled by modifying the distance metric. This formulation captures the limited and imperfect nature of real-world observations.
@@
where $r_{x}$ and $r_{y}$ are the half-lengths of the FoV rectangle along the longitudinal and lateral directions, respectively, while $\zeta$ is the forward shift in ego frame to bias the sensing region ahead of the vehicle. The rotation matrix $R(\theta)$ transforms vectors from the ego frame to the world frame.
@@
Let $O_{t}^{s}\in\{0,1\}$ be the ground-truth occupancy and $Z_{t}^{s}\in\{0,1\}$ be the observed occupancy of spot $s$ at time $t$, where $1:=$ occupied and $0:=$ vacant. Observation accuracy is the likelihood of correctly observing spot $s$, modeled using the distance-dependent probability
$p_{c}(d_{(\zeta,r)}(x,s))=P(Z_{t}^{s}=1|O_{t}^{s}=1)=P(Z_{t}^{s}=0|O_{t}^{s}=0),$ (4)
$p_{c}(d_{(\zeta,r)}(x,s))=e^{-\ln(2)\left(1+e^{-\alpha_{c}\left(d_{(\zeta,r)}(x,s)-\left(\frac{\gamma-\epsilon}{2}+\epsilon\right)\right)}\right)^{-1}},$
where $p_{c}(d)$ is the exponential of a scaled sigmoid centered at $\frac{\gamma-\epsilon}{2}+\epsilon$, such that $p_{c}(d)\to 1$ as $d$ decreases to $\epsilon$ and $p_{c}(d)$ decreases to $0.5$ as $d\to\gamma$.The probability (4) models observation confidence degrading with distance from ego vehicle, unlike prior work [23] that assumes fixed uncertainty.
@@
Given the ego vehicle state $x_{t}\in\mathbb{R}^{3}$ at time $t$, parking spot observations $Z_{t}^{s}$ as defined in Section II-C, predictions of $D$ dynamic agents $\{\{\hat{y}^{j}(k)\}_{k=t}^{t+T}\}_{j=1}^{D}$ over a horizon $T$, where $\hat{y}^{j}(k)\in\mathbb{R}^{3}$ is the position and heading of agent $j$ if it is a vehicle, and $\hat{y}^{j}(k)\in\mathbb{R}^{2}$ is the position if agent $j$ is a pedestrian; compute a reference trajectory that satisfies the vehicle dynamics (1) and steers the vehicle toward a feasible parking spot or explore the lot while avoiding static and dynamic obstacles.
Remark: We assume the availability of predicted trajectories of dynamic agents over a horizon $T$ at each planning step $t$ to solve Problem 1, while aiming to develop and integrate a trajectory prediction model in future work.
@@ II-D Proposed Approach
Fig. 3 illustrates our approach, which integrates spot occupancy estimation with planning. The perception module provides dynamic agent predictions and occupancy observations using our partial FoV model, and the occupancy estimator recursively updates to predict future spot availability. The strategy planner then computes and evaluates feasible paths, corresponding to one of the three actions—park immediately, wait near a potential spot, or explore the lot—by minimizing a cost that trades off efficiency and safety. The chosen trajectory is executed by the vehicle controller, closing the loop.
III Spot Occupancy Estimator
At each planning step $t$, the spot occupancy estimator predicts occupancy probabilities for all spots in the lot over a horizon $T$, combining current observations with predicted motion of dynamic agents. This enables the ego vehicle to plan based on anticipated occupancies rather than instantaneous observations. Probabilities are computed via a Bayes filter [28], with the prediction step modeling agent interactions and the update step incorporating FoV observations.
@@ III-A Effect of Dynamic Agents
Each agent $j$’s position at time $k$ is modeled as a Gaussian:
@@
where $\hat{y}^{j}(k)$ is the nominal prediction and $\Sigma(k)$ is the covariance. These uncertain trajectories define the probability that a spot $s\in\mathcal{P}$ is occupied by an agent $j$ at time $k$:
@@
where
$E_{d}=\big{(}(\alpha_{d}+1)(\alpha_{d}+e^{\alpha_{1}m_{d}(y^{j}(k),s)})\big{)}^{-1}$ (7)
captures spatial proximity via the Mahalanobis distance [29]:
$m_{d}(y^{j}(k),s)=(\hat{y}^{j}(k)-s)^{T}\Sigma(k)^{-1}(\hat{y}^{j}(k)-s)$
and
@@
is an alignment factor that reflects whether the agent is moving toward the spot, based on its predicted velocity $\hat{v}^{j}(k)$ and velocity uncertainty $\sigma_{v}(k)$. Together, $E_{d}$ and $A_{f}$ encode the intuition that an agent is more likely to occupy a spot if the agent is both closer and moving towards the spot, while accounting for motion uncertainty.
Uncertainty Propagation: We propagate the positional covariance at each timestep using the discrete-time model:
@@
where $R(\theta(k))$ is the rotation matrix, $\theta(k)$ is vehicle’s heading and $Q$ is a diagonal matrix representing process noise in ego frame. The initial covariance is $\Sigma(t)=\begin{bmatrix}\beta_{l}L&0\\ 0&\beta_{w}W\end{bmatrix}$, with $L$ and $W$ denoting the length and width of the vehicle, while $\beta_{l}$ and $\beta_{w}$ correspond to the respective scaling factors. If the dynamic agent is a pedestrian, then $\theta(k)=0\deg$, $\beta_{l}=\beta_{w}=1$, and we use a radius parameter to represent the size of the pedestrian instead of the length and width of the vehicle. To estimate the uncertainty in velocity, we approximate the instantaneous velocity as the finite difference $v^{j}(k)=\frac{y^{j}(k+1)-y^{j}(k)}{\Delta t}$, where $\Delta t$ is the time interval between successive time steps. The resulting covariance in velocity is
From (6), the probability that at least one of the $D$ dynamic agents will occupy spot $s$ at time $k$ is:
@@
At each step $k\in\{t,\dots,t+T\}$, we predict the prior belief $\bar{b}_{k}^{s}=P(O_{k}^{s}=1)$ using (10), and update it with the observations $Z_{k}^{s}$ to obtain the posterior belief $b_{k}^{s}=P(O_{k}^{s}=1\mid Z_{k}^{s})$}, as described in Section III-B.
@@ III-B Prediction
Prior work [23] propagates beliefs without considering the effect of dynamic agents on parking spots. We explicitly incorporate dynamic agent motion and model asymmetric interaction with spots: the ego cares about a currently occupied spot becoming vacant, while it is concerned about a vacant spot becoming occupied in the future. This distinction enables faster parking and smoother trajectories, as shown in Section V-A1. The current occupancy of a spot is determined from its observation $Z_{t}^{s}$. No occupancy observations and agent interactions are available outside the FoV.
@@
    where $q(s,k+1)=P(O_{k+1}^{s}=1|O_{k}^{s}=0)$ is computed from (10) and $\mu_{1}=P(O_{k+1}^{s}=1|O_{k}^{s}=1)$ models the departure probability of a parked vehicle using an exponential distribution [30], reflecting the memory-less nature of departures. If $\lambda_{d}$ denotes the average departure rate of a parked vehicle, the probability of departure within a time interval $\Delta t$ is
@@
    where $1-q(s,k+1)=P(O_{k+1}^{s}=0|O_{k}^{s}=1)$ is computed from (10) and $\mu_{2}=P(O_{k+1}^{s}=1|O_{k}^{s}=0)$ models the arrival probability of a vacant spot using a Poisson distribution, which commonly describes arrival processes [31]. If $\lambda_{a}$ denotes the average number of vehicles that arrive with the time interval $\Delta t$, the probability that no vehicle parks in a vacant spot is
@@
    For unobservable spots, the prediction equation depends only on the average departure and arrival rates:
@@
The prior $\bar{b}_{k+1}^{s}$ serves as the input to the update step, where it is fused with observations.